OpenStudy (anonymous):

Two boys want to use a seesaw, and they need to move the seesaw so that their weights will balance out. The formula is given by w1 • d1 = w2 • d2 where w1 = weight of the first boy, d1 = distance of the first boy from the fulcrum, w2 = weight of the second boy, and d2 = distance of the second boy from the fulcrum. Rewrite the formula to solve for d2. Show all steps in your work.

OpenStudy (welshfella):

w1 . d1 = w2 . d2 divide both sides by w2 and you have it.

OpenStudy (anonymous):

Oh ok that's not that hard at all so u understand when one thing multiplies on one side it divides on other

OpenStudy (anonymous):

So what they are saying make d2 the subject

OpenStudy (anonymous):

Which means d2 must be left alone in one side so you w2 is multiplying by d2 just divide it in the other side

OpenStudy (anonymous):

w1 x d1 = w2 x d2 1 = 4 ?? im soz im confused by this

OpenStudy (anonymous):

Where did u get 1=4

OpenStudy (anonymous):

well multiplying idek how ;-;

OpenStudy (goose.):

Do you have any numbers to plug in?

OpenStudy (anonymous):

No you won't get 1=4 you will get d2= ( w1 times d1)/ w2

OpenStudy (anonymous):

There are no numbers in question which u need to plug in

OpenStudy (welshfella):

you are give no values yor answer will be a formula with d2 as the subject

imqwerty (imqwerty):

\[w _{1}\times d _{1}=w _{2}\times d _{2}\]divide both sides by w2 u get-\[\frac{ (w _{1} \times d _{1}) }{ w _{2} }=d _{2}\]

OpenStudy (welshfella):

exactly - that's what i said to do in my first post

OpenStudy (anonymous):

so then its just w2 = d2 thats the final answer

OpenStudy (anonymous):

?*

OpenStudy (welshfella):

No the answer is the formula that imqwerty gave in his post

OpenStudy (welshfella):